Session: May 1999 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Sat Jun 12 10:20:58 METDST 1999) Success
Experiment OK. All scans recorded.

Noto (C.Trigilio at Mon Jun 14 08:19:46 METDST 1999) Success

Effelsberg (A. Kraus at Mon Jun 14 09:04:22 METDST 1999) Minor failures
Problems with the position of tape head one occured between UT 0642 - 0700 and UT 0720 - 0723.

Onsala (A. Polatidis at Mon Jun 14 16:44:05 METDST 1999) Success

Torun (K. Borkowski at Tue Jun 15 14:47:18 METDST 1999) Success
No known problems.

Jodrell (Alastair Gunn at Fri Jun 18 10:58:38 METDST 1999) Success
No known problems.

Cambridge (Alastair Gunn at Fri Jun 18 10:59:04 METDST 1999) Minor failures
Bad data for first few minutes due to difficulty changing LO values. No other known problems.

Westerbork (Tony Foley at Fri Jun 18 16:09:48 METDST 1999) Minor failures
Used 13 telescopes in the array. On a brief parity check the error rates seemed high - Socorro please check.

Calibration (Jean\-Francois Desmurs at Mon Jun 21 16:13:50 METDST 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may99 directory under the name: es030a.antab. Read the comments in the file to get further information. (Rem : Cm no Tsys values.)